OpenStudy (anonymous):

When solving negative one over three.(x − 15) = −4, what is the correct sequence of operations? Multiply each side by negative one over three , subtract 15 from each side Multiply each side by −3, add 15 to each side Multiply each side by negative one over three , add 15 to each side Multiply each side by 3, subtract 15 from each side

OpenStudy (texaschic101):

-1/3(x - 15) = -4--- multiply each side by -3 (-3/1)(-1/3) (x - 15) = (-3)(-4) (x - 15) = 12 --- add 15 to both sides x = 12 + 15 x = 27

OpenStudy (anonymous):

so b

OpenStudy (cutebunny12):

is b the right answer?